# PiShield: A NeSy Framework for Learning with Requirements

### v1

@@ Abstract
Deep learning models have shown their strengths in various application domains, however, they often struggle to meet safety requirements for their outputs. In this paper, we introduce PiShield, the first framework ever allowing for the integration of the requirements into the neural networks’ topology. PiShield guarantees compliance with these requirements, regardless of input. Additionally, it allows for integrating requirements both at inference and/or training time, depending on the practitioners’ needs. Given the widespread application of deep learning, there is a growing need for frameworks allowing for the integration of the requirements across various domains. Here, we explore three application scenarios: functional genomics, autonomous driving, and tabular data generation.
@@ 1 Introduction
In this paper, we propose PiShield [^1] Code: ${}^{,}$ [^2] Website: https://sites.google.com/view/pishield, a PyTorch-based framework allowing for seamlessly integrating domain requirements into neural networks by means of new PyTorch layers that can be built on top of any neural network. These layers, which we call Shield Layers, adhere to the principles outlined in our most recent works Stoian et al. (2024); Giunchiglia et al. (2024) and guarantee the satisfaction of the requirements regardless of the input.
@@ 3 Example Scenarios
Generalizing the above problem, we have the task of assigning sets of labels to instances of various types, including images, documents etc. Given available domain knowledge, constraints are placed on the possible outputs, eliminating impossible (i.e., contradicting background knowledge) outcomes. In our recent work Giunchiglia et al. (2024), we considered the task of multi-label classification for an autonomous driving scenario using the ROAD-R dataset Giunchiglia et al. (2023), the first real-world dataset for autonomous driving with manually annotated logical constraints. ROAD-R is build upon the road event detection dataset ROAD Singh et al. (2023) by adding 243 propositional logic requirements written in CNF. We showed that applying Shield Layers on the outputs is not only able to guarantee that these requirements are satisfied, but also results in increased performance, as seen in Table 2 of Giunchiglia et al. (2024). Figure 2 shows examples of how using PiShield can impact the predictions’ quality, which is intrinsically coupled with the safety guarantees.
@@ 4 Conclusions
In this paper, we introduced PiShield, the first framework that allows for injecting requirements into neural networks’ topology by building Shield Layers, which correct the outputs so that they are guaranteed to satisfy the rules. PiShield can be applied (i) during training by building Shield Layers into the neural networks’ architecture, or (ii) at inference time by correcting the neural networks’ outputs as a post-processing step. We envision our framework will be of use to practitioners working on real-world applications where domain knowledge can be expressed as propositional or linear constraints.

### v2

# PiShield: A PyTorch Package for Learning with Requirements

@@ Abstract
Deep learning models have shown their strengths in various application domains, however, they often struggle to meet safety requirements for their outputs. In this paper, we introduce PiShield, the first package ever allowing for the integration of the requirements into the neural networks’ topology. PiShield guarantees compliance with these requirements, regardless of input. Additionally, it allows for integrating requirements both at inference and/or training time, depending on the practitioners’ needs. Given the widespread application of deep learning, there is a growing need for frameworks allowing for the integration of the requirements across various domains. Here, we explore three application scenarios: functional genomics, autonomous driving, and tabular data generation.
@@ 1 Introduction
In this paper, we propose PiShield [^1] Code: https://github.com/mihaela-stoian/PiShield, [^2] Website: https://sites.google.com/view/pishield, a PyTorch-based package allowing for seamlessly integrating domain requirements into neural networks by means of new PyTorch layers that can be built on top of any neural network. These layers, which we call Shield Layers, adhere to the principles outlined in our most recent works Stoian et al. (2024); Giunchiglia et al. (2024), advocating for a more requirements-driven machine learning as in Giunchiglia et al. (2023a), and guarantee the satisfaction of the requirements regardless of the input.
@@ 3 Example Scenarios
Generalizing the above problem, we have the task of assigning sets of labels to instances of various types, including images, documents etc. Given available domain knowledge, constraints are placed on the possible outputs, eliminating impossible (i.e., contradicting background knowledge) outcomes. In our recent work Giunchiglia et al. (2024), we considered the task of multi-label classification for an autonomous driving scenario using the ROAD-R dataset Giunchiglia et al. (2023b), the first real-world dataset for autonomous driving with manually annotated logical constraints. ROAD-R is build upon the road event detection dataset ROAD Singh et al. (2023) by adding 243 propositional logic requirements written in CNF. We showed that applying Shield Layers on the outputs is not only able to guarantee that these requirements are satisfied, but also results in increased performance, as seen in Table 2 of Giunchiglia et al. (2024). Figure 2 shows examples of how using PiShield can impact the predictions’ quality, which is intrinsically coupled with the safety guarantees.
@@ 4 Conclusions
In this paper, we introduced PiShield, the first package that allows for injecting requirements into neural networks’ topology by building Shield Layers, which correct the outputs so that they are guaranteed to satisfy the rules. PiShield can be applied (i) during training by building Shield Layers into the neural networks’ architecture, or (ii) at inference time by correcting the neural networks’ outputs as a post-processing step. We envision our package will be of use to practitioners working on real-world applications where domain knowledge can be expressed as propositional or linear constraints.